Local Corp. has an EPS of $3.50 and an earnings growth rate of 8%. If the benchmark P/E ratio is 17, what is the target stock price 3 years from now?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

EPS would be=3.5*(1.08)^3

=$4.408992

P/E ratio=stock price/EPS

Hence target price=4.408992*17

=$74.95(Approx).